Suppose the American Kennel Club estimates that 36% of households in U.S. have a Pembroke Welsh Corgi. A random sample of 400 households is taken. What is the probability the sampling error is 5% or less?
Answer #1

Ans:

n=400

z(0.31)=(0.31-0.36)/SQRT(0.36*(1-0.36)/400)=-2.083

z(0.41)=(0.41-0.36)/SQRT(0.36*(1-0.36)/400)=2.083

P(0.31<p-hat<0.41)=P(-2.083<z<2.083)=P(z<2.083)-P(z<-2.083)

=0.9814-0.0186=0.9628